OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42.

OpenStudy (phebe):

\[4x-2x-2=-42\]

OpenStudy (anonymous):

4x-2x=2x okay now put at the place of both these two points only 2x

OpenStudy (anonymous):

x= -10

OpenStudy (phebe):

do like terms first

OpenStudy (anonymous):

-10 is x

OpenStudy (phebe):

yeaa @Jello_Soccer is right its -10

OpenStudy (anonymous):

2x-2=-42 2x=-42+2 2x=-40 x=40/2 X=-20 is the correct answers.

OpenStudy (anonymous):

Thank you rio ^.^

OpenStudy (anonymous):

im wrong 0.0

OpenStudy (anonymous):

oh well

OpenStudy (phebe):

ohh shoot i was wrong tooo well thanx for explaing then @rio96

OpenStudy (anonymous):

yea ty rio

OpenStudy (anonymous):

Its ok thanks all of you (: